if third term of gp is 2/9 and sum of infinite terms is 3/2 then find the series

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
ar2 = 2/9a/(1-r) = 3/2dividing the equtions r2(1-r)= 4/27=> r=2/3and a=9/4*(2/9)a=1/2so series is ½, 1/3, 2/9, 4/27...